# Certified algorithms for equilibrium states of local quantum Hamiltonians

### v2

@@ 3.5 Commuting Hamiltonians and the high-temperature regime
In this section we focus on commuting Hamiltonians $H$ which satisfy $[h_{X},h_{Y}]=0$ for any $X,Y\subset\Gamma$. For concreteness, we choose to focus on $H$ translation-invariant on $\Gamma=\mathbb{Z}^{D}$. For such Hamiltonians one can define a convex relaxation for the set of thermal equilibrium states directly from the KMS equality condition. Indeed, in this case if $b\in\mathcal{A}_{\Lambda}\subset\mathcal{A}_{\text{loc}}$ where $\Lambda$ is finite, then
@@
[…]
@@
We are now ready to prove exponential convergence of the convex optimization hierarchy to $\langle O\rangle_{\beta}$ for $0\leq\beta<\beta_{1}$. We start with the 1D case. Our main tool will be the decay of correlations and local indistinguishability of Gibbs states for one-dimensional chains [56]. Without loss of generality let us assume a 2-local interaction; the general case follows from blocking the sites first. Let $\omega\in\mathcal{G}_{\Lambda_{\ell},\beta}^{\text{comm}}$ and consider the perturbation $\omega^{W}$ from Lemma 3.10. The surface interaction term $W$ has support on the sites $-(\ell+1),-\ell,\ell,\ell+1$ and $\|\exp(\pm\beta W)\|\leq\exp(2\beta\|h\|)$ is bounded by a constant. We first apply the decay of correlations to bound the effect of the perturbation. From Lemma 3.10 we have
@@ 4 Numerical experiments
When $g=0$, the model is a simple one-dimensional Ising model, and there are two translation-invariant ground states, namely the all-up and all-down states. This means that at $g=0$ the magnetization can take any value between $-1$ and $+1$. For general $g\geq 0$, it has been shown that [60, Eq. (3.12a)]
@@
[…]
@@
[…]
@@
[…]
@@ Acknowledgments
We thank Daniel Stilck França for helpful discussions. HF acknowledges funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee EP/X032051/1. OF acknowledges funding by the European Research Council (ERC Grant AlgoQIP, Agreement No. 851716) as well as by the European Union’s Horizon 2020 within the QuantERA II Programme under Grant VERIqTAS Agreement No 101017733. SOS acknowledges support from the UK Engineering and Physical Sciences Research Council (EPSRC) under grant number EP/W524141/1.
@@
[…]
@@
We see from the second part that $[\mu(1)]_{kl}=\omega(a_{k}^{*}a_{l})=A_{kl}$ and $\nu(1)=B$. Furthermore, we show that $d\nu(p)=e^{p\beta}d\mu(p)$ following the analogous statement in [45, Proposition 5.3.14]. We do this by computing for a compactly supported infinitely differentiable function $\hat{f}$
[…]
@@
[…]

### v3

@@ 3.5 Commuting Hamiltonians and the high-temperature regime
In this section we focus on commuting Hamiltonians $H$ which satisfy $h_{X},h_{Y}]=0$ for any $X,Y\subset\Gamma$. For concreteness, we choose to focus on $H$ translation-invariant on $\Gamma=\mathbb{Z}^{D}$. For such Hamiltonians one can define a convex relaxation for the set of thermal equilibrium states directly from the KMS equality condition. Indeed, in this case if $b\in\mathcal{A}_{\Lambda}\subset\mathcal{A}_{\text{loc}}$ where $\Lambda$ is finite, then
@@
$\mathcal{G}_{\Lambda,\beta}^{\text{comm}}=\Bigl{\{}\omega:\mathcal{A}_{\bar{\Lambda}}\to\mathbb{C}\text{ s.t. }$ $\omega(1)=1,\omega(a^{*}a)\geq 0\;\;\forall a\in\mathcal{A}_{\bar{\Lambda}},$ (67)
@@
We are now ready to prove exponential convergence of the convex optimization hierarchy to $\langle O\rangle_{\beta}$ for $0\leq\beta<\beta_{1}$. We start with the 1D case. Our main tool will be the decay of correlations and local indistinguishability of Gibbs states for one-dimensional chains [56]. Without loss of generality let us assume a 2-local interaction; the general case follows from blocking the sites first. Let $\omega\in\mathcal{G}_{\Lambda_{\ell},\beta}^{\text{comm}}$ and consider the perturbation $\omega^{W}$ from Lemma 3.10. The surface interaction term $W$ has support on the sites $(\ell+1),-\ell,\ell,\ell+1$ and $\|\exp(\pm\beta W)\|\leq\exp(2\beta\|h\|)$ is bounded by a constant. We first apply the decay of correlations to bound the effect of the perturbation. From Lemma 3.10 we have
@@ 4 Numerical experiments
When $g=0$, the model is a simple one-dimensional Ising model, and there are two translation-invariant ground states, namely the all-up and all-down states. This means that at $g=0$ the magnetization can take any value between 1-1 and +1+1. For general $g\geq 0$, it has been shown that [60, Eq. (3.12a)]
@@
00.20.20.40.40.60.60.80.8111.21.21.41.41.61.61-10.5-0.500.50.511ggMagnetization $\langle Z_{0}\rangle$ at $\beta=\inftyL$=4L=4L=5L=5
00.20.20.40.40.60.60.80.8111.21.21.41.41.61.61-10.5-0.500.50.511ggMagnetization $\langle Z_{0}\rangle$ at $\beta=1$L=4L=4L=5L=5
@@
$\langle Z_{0}Z_{1}\rangle_{g,\beta}=\frac{1}{\pi}\int_{0}^{\pi}\frac{1+g\cos k}{\sqrt{1+2g\cos k+g^{2}}}\tanh(J\beta\sqrt{1+2g\cos k+g^{2}})dk.$ (92)
@@
0.50.5111.51.500.50.511ggCorrelation $\langle Z_{0}Z_{1}\rangle$ at $\beta=\inftyL$=4L=4L=5L=5
0.50.5111.51.50.20.20.40.40.60.60.80.811ggCorrelation $\langle Z_{0}Z_{1}\rangle$ at $\beta=1$L=4L=4L=5L=5
@@ Acknowledgments
We thank Daniel Stilck França for helpful discussions. We thank David Pérez-García for pointing out the necessity of using subnets in the proof of Theorem 3.6. HF acknowledges funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee EP/X032051/1. OF acknowledges funding by the European Research Council (ERC Grant AlgoQIP, Agreement No. 851716) as well as by the European Union’s Horizon 2020 within the QuantERA II Programme under Grant VERIqTAS Agreement No 101017733. SOS acknowledges support from the UK Engineering and Physical Sciences Research Council (EPSRC) under grant number EP/W524141/1.
@@
$[\mu(\hat{f})]_{kl}$ $=(2\pi)^{-1/2}\int_{-\infty}^{\infty}dtf(t)\omega(a_{k}^{*}\alpha_{t}(a_{l}))=\int_{-\infty}^{\infty}\langle\pi(a_{k})\Omega,dE(q)\pi(a_{l})\Omega\rangle\hat{f}(q)dq$ (94)
$[\nu(\hat{f})]_{kl}$ $=(2\pi)^{-1/2}\int_{-\infty}^{\infty}dtf(t)\omega(\alpha_{t}(a_{l})a_{k}^{*})=\int_{-\infty}^{\infty}\langle\pi(a_{k})\Omega,dE(-q)\pi(a_{l})\Omega\rangle\hat{f}(q)dq$ (95)
@@
We see from the second part that $\mu(1)]_{kl}=\omega(a_{k}^{*}a_{l})=A_{kl}$ and $\nu(1)=B$. Furthermore, we show that $d\nu(p)=e^{p\beta}d\mu(p)$ following the analogous statement in [45, Proposition 5.3.14]. We do this by computing for a compactly supported infinitely differentiable function $\hat{f}$
$[\mu(\hat{f})]_{kl}$ $=\int_{\infty}^{\infty}dtf(t)\omega(a_{k}^{*}\alpha_{t}(a_{l}))$ (97)
$=\int_{\infty}^{\infty}dtf(t+i\beta)\omega(\alpha_{t}(a_{l})a_{k}^{*})=[\nu(k_{\beta}\hat{f})]_{kl}$ (98)
@@
$D_{op}(A\|B)$ $=D_{op}\left(\int 1d\mu(p)\|\int 1d\nu(p)\right)$ (99)
$=D_{op}\left(\int\tilde{A}(p)d\gamma(p)\|\int e^{p\beta}\tilde{A}(p)d\gamma(p)\right)$
$\preceq\int D_{op}(\tilde{A}(p)\|e^{p\beta}\tilde{A}(p))d\gamma(p)$